sherry akey invested some money at 8% and $100 more than that at 6%. her total annual interest was $48. how much did she invest at each rate?
Answer by stanbon(75887)

sherry akey invested some money at 8% and $100 more than that at 6%. her total annual interest was $48. how much did she invest at each rate?
------
Equation:
int + int = int
0.08x + 0.06(x+100) = 48
-----
8x + 6x + 600 = 4800
14x = 4200
x = $300 (amt. invested at 8%
---
x + 100 = $400 (amt. invested at 6%)
===================
Cheers,
Stan H.
--------------------